# The optimal hypercontractive constants for $\mathbb{Z}_{3}$

### v1

@@ Abstract.
We resolve a folklore problem for the optimal hypercontractive constant of the cyclic group $\mathbb{Z}_{3}$ for all $1<p<q<\infty$. Precisely, the optimal hypercontractive constant is given by
@@
One feature of our formalism is that the system simultaneously determines the optimal hypercontractive constant and a nontrivial extremizer. As a consequence of our main result, for rational $p,q\in\mathbb{Q}$, the constants $r_{p,q}(\mathbb{Z}_{3})$ are algebraic numbers whose minimal polynomials are generally rather complicated, with non-solvable Galois groups and therefore no explicit radical expressions.
@@
Hypercontractivity, hypercontractive constant, nontrivial extremizers, variational principle, blowup analysis
@@ 1.1. Main results
Given an integer $n\geq 2$, let $\mathbb{Z}_{n}=\mathbb{Z}/n\mathbb{Z}$ denote the finite cyclic group equipped with the Haar measure $\mathfrak{m}$ normalized by $\mathfrak{m}(\mathbb{Z}_{n})=1$. Throughout the paper, let $\chi\in\widehat{\mathbb{Z}_{n}}$ be the character with $\chi(1)=e^{i2\pi/n}$. For any $r\in[0,1]$ and any $p>1$, define the dilation operator on $L^{p}(\mathbb{Z}_{n})=L^{p}(\mathbb{Z}_{n},\mathfrak{m})$ as
$T_{r}:\sum_{k=0}^{n-1}a_{k}\chi^{k}\mapsto\sum_{k=0}^{n-1}a_{k}r^{d(k)}\chi^{k},$
where $a_{k}\in\mathbb{C}$ for all $0\leq k\leq n-1$ and $d(k)$ denotes the graph distance given by $d(k)=\min\{k,n-k\}$ for $0\leq k\leq n-1$.
@@
$\left\|T_{r}f\right\|_{q}\leq\|f\|_{p}\quad\text{for all $0\leq r\leq r_{p,q}(\mathbb{Z}_{n})$ and all $f\in L^{p}(\mathbb{Z}_{n})$.}$
A non-constant function $f$ for which the hypercontractive inequality becomes an equality when $r=r_{p,q}(\mathbb{Z}_{n})$ is called a non-trivial $(p,q)$-extremizer for $\mathbb{Z}_{n}$ or simply a non-trivial extremizer.
@@
And, it has been established that the hypercontractive constant coincides with that of $\mathbb{Z}_{2}$ for many other cases on finite cyclic groups.
@@
However, the hypercontractivity of $\mathbb{Z}_{3}$ differs drastically from all the established cases and some partial results on $r_{p,q}(\mathbb{Z}_{3})$ were obtained by Andersson [And02], Diaconis and Saloff-Coste [DSC96]. And Wolff proved in [Wol07, Corollary 3.1] that
$r_{2,q}(\mathbb{Z}_{3})=\sqrt{2(4^{1/q}-1)/(4-4^{1/q})}.$
Wolff’s result is based on Latała and Oleszkiewicz [LaO00, Ole03], where the condition $p=2$ is crucially used. More details and applications of hypercontractive inequalities on discrete space can be found in [And99, CLSC08, O’D14, JPPR15, IN22, KLLM24, SVZ24] and references therein.
@@
This inspires us to move away from seeking an explicit closed-form expression for $r_{p,q}(\mathbb{Z}_{3})$, which appears unlikely to exist, and instead to analyze its implicit functional structure. Through this shift in perspective, we succeed in determining $r_{p,q}(\mathbb{Z}_{3})$ for all $1<p<q<\infty$, thereby resolving the problem.
@@
Moreover, the function $f=1+\frac{1-x}{1+2x}(\chi+\overline{\chi})$ is a nontrivial extremizer.
In seeking the optimal constant $r_{p,q}(\mathbb{Z}_{3})$, we unexpectedly discover the existence of nontrivial $(p,q)$-extremizers, which is in sharp contrast to the classical situation. The existence of such non-trivial extremizer seems to be related to the strict inequality
$r_{p,q}(\mathbb{Z}_{3})<\sqrt{(p-1)/(q-1)}$
and it naturally allows us to obtain an equation-system via the simplest variational principle (that is, the Fermat principle). However, the uniqueness of solution to the corresponding equation-system is highly nontrivial: indeed, inspired by the theory of Whitney pleats [AGV85, §1], we are led to analyze the intersection properties of a curve-family associated to (1.2) by using symmetrization and blowup techniques.
@@
The formula (1.1) for $r_{p,q}(\mathbb{Z}_{3})$ is exactly the cross ratio (see, e.g., [Bea95, §4.4] for its definition) of x,y,1/2,1x,y,-1/2,1:
(1.3) $r_{p,q}(\mathbb{Z}_{3})=(x,y;-1/2,1):=\frac{(x+1/2)(y-1)}{(x-1)(y+1/2)}.$
@@
where $p^{*}=p/(p-1),q^{*}=q/(q-1)$ are the corresponding conjugate exponents. On the other hand, the equation-system (1.2) also processes a symmetry of duality and admits a self-dual equivalent formulation as follows. For any $1<p<\infty$ and $x\in(0,1)$, define
$\ell(p,x):=\frac{1}{1+2x}\Big(\frac{1+2x^{p}}{3}\Big)^{\frac{1}{p}}.$
@@
The equation-system (1.2) has the following equivalent formulation:
@@
And the cross ratio (x,y;1/2,1)(x,y;-1/2,1) for the solution $(x,y)$ of the equation-system (1.2) is also invariant under the above transform:
@@
In general, $r_{p,q}(\mathbb{Z}_{3})$ rarely admits an explicit expression; the following corollary presents the rare instances that we conjecture to be the only ones. Note that, the expression $r_{2,p^{*}}(\mathbb{Z}_{3})=r_{p,2}(\mathbb{Z}_{3})$ was obtained by Wolff [Wol07], while the expression for $r_{p,p^{*}}(\mathbb{Z}_{3})$ appears to be new in the literature, and one may easily see from the proof in §C.2 that the explicit expression for $r_{p,p^{*}}(\mathbb{Z}_{3})$ is closely related to the symmetry (1.5).
@@
Note that $r_{p,q}(\mathbb{Z}_{2})=r_{p,s}(\mathbb{Z}_{2})r_{s,q}(\mathbb{Z}_{2})$ for all $1<p\leq s\leq q<\infty$. This multiplicative relation does not generally hold for $r_{p,q}(\mathbb{Z}_{3})$. This fact seems to be a common understanding and can be seen from the numerical solutions of the system (1.2) in the Figure 1. The lack of this general multiplicative relation seems another obstacle in expecting an explicit formula for $r_{p,q}(\mathbb{Z}_{3})$.
Figure 1. Nonmultiplicative relationship
@@
Although the derivation of Corollary 1.4 from Theorem 1.1 seems rather elementary in algebraic geometry, for the reader’s convenience we provide a short proof in the appendix D that does not require any background in the subject.
Remark.
For $p,q\in\mathbb{Q}$, the minimal polynomial of $r_{p,q}(\mathbb{Z}_{3})$ may be written explicitly. Indeed, in view of Theorem 1.1 or the equivalent equation-system (4.2) concerning directly $r_{p,q}(\mathbb{Z}_{3})$, one can use the standard method of resultant (see the proof of Corollary 1.4 in the appendix D for a quick explanation) to derive an integer-coefficient polynomial (and then an irreducible one) having $r_{p,q}(\mathbb{Z}_{3})$ as a root. For instance, using computer algebra systems (e.g., Mathematica, Matlab), we obtain the minimal polynomial of $r_{3,6}(\mathbb{Z}_{3})$:

### v2

# arXiv:2602.17248v2[math.FA] 01 Mar 2026

@@ Abstract.
We resolve a folklore problem of determining the optimal hypercontractive constants $r_{p,q}(\mathbb{Z}_{3})$ for the cyclic group $\mathbb{Z}_{3}$ for all $1<p<q<\infty$. More precisely, we have
@@
Consequently, for rational $p,q\in\mathbb{Q}$, the constants $r_{p,q}(\mathbb{Z}_{3})$ are algebraic numbers which generally admit no radical expressions, since their often rather complicated minimal polynomials may have non‑solvable Galois groups.
Our formalism relies on a key observation: the existence of nontrivial critical extremizers. This approach can also be adapted to resolve a long-standing open problem—determining all optimal $(p,q)$-hypercontractive constants for biased Bernoulli random variables, which are closely related to noise operators. Several noteworthy phenomena emerge from numerical simulations: the monotonicity of the hypercontractive constants in the parameters, and the appearance of intriguing limit shapes. These phenomena merit further investigation.
@@
Hypercontractivity, optimal hypercontractive constant, nontrivial critical extremizers, biased Bernoulli random variables, blowup analysis
@@ 1.1. Main results
Given an integer $n\geq 2$, let $\mathbb{Z}_{n}=\mathbb{Z}/n\mathbb{Z}$ denote the finite cyclic group equipped with the Haar probability measure $\mathfrak{m}$ and the graph distance $d(k)=\min\{k,n-k\}$ for $0\leq k\leq n-1$. Throughout the whole paper, let $\chi\in\widehat{\mathbb{Z}_{n}}$ be the unique character determined by $\chi(1)=e^{i2\pi/n}$. For any $r\in[0,1]$ and any $p>1$, define the dilation operator on $L^{p}(\mathbb{Z}_{n})=L^{p}(\mathbb{Z}_{n},\mathfrak{m})$ as
$T_{r}:\sum_{k=0}^{n-1}a_{k}\chi^{k}\mapsto\sum_{k=0}^{n-1}a_{k}r^{d(k)}\chi^{k},\quad\text{with $a_{k}\in\mathbb{C}$ for all $0\leq k\leq n-1$.}$
@@
(1.1) $\left\|T_{r}f\right\|_{q}\leq\|f\|_{p}\quad\text{for all $0\leq r\leq r_{p,q}(\mathbb{Z}_{n})$ and all $f\in L^{p}(\mathbb{Z}_{n})$.}$
@@
And, it is proved that this hypercontractive constant is shared by many other finite cyclic groups:
@@
However, the hypercontractivity of $\mathbb{Z}_{3}$ is drastically different and partial results were obtained by Andersson [And02], Diaconis and Saloff-Coste [DSC96]. The argument of Wolff [Wol07] implies
(1.2) $r_{2,q}(\mathbb{Z}_{3})=\sqrt{2(4^{1/q}-1)/(4-4^{1/q})}.$
Wolff’s result is based on Latała and Oleszkiewicz [LaO00, Ole03], where the condition $p=2$ is crucially used. Beyond their own interest in analysis, hypercontractive inequalities on $\mathbb{Z}_{3}$ play roles in the study of the traffic light problem and subsets of independent sets in product graphs, see [ADFS04], [DFR08] and [Fri23]. More details and applications of hypercontractive inequalities on discrete space can be found in [And99, CLSC08, O’D14, JPPR15, IN22, KLLM24, SVZ24]. One can also consult [Jan83, Wei80, Gro99, Gro02, EGSC18, FI21, Kul22, Mel23] for the studies and applications of the hypercontractive and related log-Sobolev inequalities on complex manifolds.
@@
This inspires us to move away from seeking an explicit closed-form expression for $r_{p,q}(\mathbb{Z}_{3})$, which appears unlikely to exist, and instead to analyze its implicit functional structure. Through this shift in perspective, we resolving the problem by determining $r_{p,q}(\mathbb{Z}_{3})$ for all $1<p<q<\infty$.
Our work relies crucially on the following concept of nontrivial critical extremizers.
Definition.
A non-constant function $f$ on $\mathbb{Z}_{n}$ (if it exists) for which the hypercontractive inequality (1.1) becomes an equality at the critical $r=r_{p,q}(\mathbb{Z}_{n})$ is called a nontrivial $(p,q)$-critical extremizer for $\mathbb{Z}_{n}$ or simply a nontrivial critical extremizer. For any given $r>r_{p,q}(\mathbb{Z}_{n})$, an $r$-supercritical $(p,q)$-extremizer or simply supercritical extremizer (which is automatically non-constant) is defined as a function $f$ on $\mathbb{Z}_{n}$ satisfying
$1<\|T_{r}\|_{L^{p}(\mathbb{Z}_{n})\to L^{q}(\mathbb{Z}_{n})}=\|T_{r}f\|_{q}/\|f\|_{p}.$
The supercritical extremizers always exist. Indeed, since $L^{p}(\mathbb{Z}_{n})$ is of finite dimension, the standard compactness argument implies that, for any $r>r_{p,q}(\mathbb{Z}_{n})$, the $r$-supercritical extremizers always exist. In the particular case of $\mathbb{Z}_{3}$, Wolff [Wol07, Theorem 3.1] provides a remarkable qualitative description of supercritical extremizers: they must take exactly two values! This description of supercritical extremizers combined with Latała and Oleszkiewicz [LaO00, Ole03] leads to the explicit expression (1.2) of $r_{2,q}(\mathbb{Z}_{3})$. Indeed, Wolff [Wol07, Theorem 3.1] proved that, for more general operators on a finite discrete probability space, any supercritical extremizer must exhibit a two-valued feature (see §1.2 for a brief introduction). Wolff’s qualitative description for supercritical extremizers has proven to be highly influential; see, e.g., [DFR08, Mos10, MOO10, FF14, HHM18].
On the other hand, the existence of nontrivial critical extremizers is not guaranteed (although Wolff’s argument implies that all nontrivial extremizers, once exist, must be exactly two-valued). For instance, it is not hard to show that $\mathbb{Z}_{2}$ does not admit nontrivial critical extremizers for any $1<p<q<\infty$. In sharp contrast to the classical situation for $\mathbb{Z}_{2}$, partially owing to the strict inequality (see Lemma 3.1)
$r_{p,q}(\mathbb{Z}_{3})<\sqrt{(p-1)/(q-1)},$
we unexpectedly discover the existence of nontrivial $(p,q)$-critical extremizers.
Lemma 1.1.
For any $1<p<q<\infty$, there exists a nontrivial $(p,q)$-critical extremizer for $\mathbb{Z}_{3}$.
Lemma 1.1 will be proved in Proposition 4.1 in the cases $1<p<q<2$ or $1<p\leq 2<q<\infty$, while the remaining case will be proved in §5.2.
Conjecture.
One may conjecture that, for $n\geq 2$, the equality $r_{p,q}(\mathbb{Z}_{n})=\sqrt{(p-1)/(q-1)}$ holds if and only if $\mathbb{Z}_{n}$ does not admit nontrivial $(p,q)$-critical extremizer.
The existence of nontrivial critical extremizers for $\mathbb{Z}_{3}$ plays a key role in this work. Specifically, any such extremizer $f$ carries essential information about the constant $r_{p,q}(\mathbb{Z}_{3})$. Through the simplest variational principle (i.e., Fermat’s principle), this enables us to derive a system of equations involving both the Fourier coefficients of $f$ and the target constant $r_{p,q}(\mathbb{Z}_{3})$. However, for obtaining a tractable equation-system, we need to further reduce the space for candidate nontrivial critical extremizers, see §1.4.1 and §3 for the details. Note that the application of ideas from variational principles to the hypercontractive inequality is not new. The analysis of Lagrange multipliers, in particular, have been skillfully applied in works such as [Wol07], [DFR08] and [Yao25].
Our first main result is
@@
Moreover, the function $f=1+\frac{1-x}{1+2x}(\chi+\overline{\chi})$ is a nontrivial critical extremizer.
The uniqueness of the solution in $(0,1)^{2}$ to the system (1.4) is a highly nontrivial fact (which we suspect may be related to some deep result in real algebraic geometry, at least when $p,q$ are rational). For proving this uniqueness, inspired by the theory of Whitney pleats (see [AGV85, §1]), we are led to analyze the intersection properties of a curve-family associated with (1.4) by employing symmetrization and blowup techniques.
Since the proof of the uniqueness statement of Theorem 1.2 is considerably more involved than the derivation of the system (1.4) itself, here we provide an alternative description of $r_{p,q}(\mathbb{Z}_{3})$ that avoids relying on this uniqueness. Specifically, by the absence of nontrivial subcritical extremizers for $\mathbb{Z}_{3}$ (see the remark below) and the argument in the establishing (1.4), one may obtain
[…]
Even without the uniqueness result from Theorem 1.2, the system (1.4) is already quite useful, as it enables numerical computation of $r_{p,q}(\mathbb{Z}_{3})$. However, our proof of uniqueness carries both theoretical and practical significance: we observed that directly solving (1.4) numerically is highly unstable, whereas employing the symmetrization and blowup techniques developed in the uniqueness proof yields an efficient and stable numerical simulation of $r_{p,q}(\mathbb{Z}_{3})$.
@@
For any $1<p<q<\infty$, one can define subcritical extremizers for $0<r<r_{p,q}(\mathbb{Z}_{n})$. However, there is no nontrivial subcritical extremizer for $\mathbb{Z}_{n}$, since $T_{r}=T_{r/r_{p,q}(\mathbb{Z}_{n})}\circ T_{r_{p,q}(\mathbb{Z}_{n})}$ and only the constant functions can have $T_{r/r_{p,q}(\mathbb{Z}_{n})}$-invariant $L^{q}$-norms.
Remark.
As a consequence of the uniqueness of the solution to (1.4), for any $1<p<q<\infty$, one may prove that, up to multiplication by a non-zero constant, there exists a unique nontrivial critical extremizer for $\mathbb{Z}_{3}$ (which however will vary for different pairs $(p,q)$).
Remark.
The formula (1.3) for $r_{p,q}(\mathbb{Z}_{3})$ is exactly a cross ratio (see, e.g., [Bea95, §4.4]):
$r_{p,q}(\mathbb{Z}_{3})=(x,y;-1/2,1):=\frac{(x+1/2)(y-1)}{(x-1)(y+1/2)}.$
@@
Remark.
In [Bor82], Borell established a reverse hypercontractive inequality on $\mathbb{Z}_{2}$. Recently, Mossel, Oleszkiewicz, and Sen [MOS13] showed that reverse hypercontractive inequalities follow from standard hypercontractive inequalities or the modified log-Sobolev inequality. In particular, they obtained that Borell’s reverse hypercontractive inequalities hold uniformly for all probability spaces. It seems that the method in this paper can be adapted to obtain the optimal constant of the reverse hypercontractivity on $\mathbb{Z}_{3}$. Indeed, one can show (with a little effort) that the equation-system (1.4) has a unique solution in $(0,1)^{2}$ when $0<q<p<1$. Thus, the optimal reverse hypercontractive constants also have the form (1.3).
@@
where $p^{*}=p/(p-1),q^{*}=q/(q-1)$ are the conjugate exponents. Meanwhile, the equation-system (1.4) also possesses a symmetry of duality. For any $1<p<\infty$, define
$\ell(p,x):=\frac{1}{1+2x}\Big(\frac{1+2x^{p}}{3}\Big)^{\frac{1}{p}},\quad x\in(0,1).$
@@
The equation-system (1.4) has the following self-dual equivalent formulation:
@@
This transform also leaves invariant the cross ratio (x,y;1/2,1)(x,y;-1/2,1) for the solution $(x,y)$ of (1.4):
@@
In general, $r_{p,q}(\mathbb{Z}_{3})$ rarely admits an explicit expression; the following corollary presents the rare instances that we conjecture to be the only ones. Note that, $r_{2,p^{*}}(\mathbb{Z}_{3})=r_{p,2}(\mathbb{Z}_{3})$ is due to Wolff [Wol07], while $r_{p,p^{*}}(\mathbb{Z}_{3})$ appears to be new in the literature. One shall see in §C.2 that these expressions are related to the symmetry (1.7).
@@
While $r_{p,q}(\mathbb{Z}_{2})=r_{p,s}(\mathbb{Z}_{2})r_{s,q}(\mathbb{Z}_{2})$ for all $1<p\leq s\leq q<\infty$, it seems to be a common understanding that the multiplicative relation does not generally hold for $r_{p,q}(\mathbb{Z}_{3})$. See also the numerical solutions of the system (1.4) in Figure  1. The non-multiplicative relation seems another obstacle in expecting an explicit formula for $r_{p,q}(\mathbb{Z}_{3})$.
Figure 1. Non-multiplicative relationship
@@
The derivation of Corollary 1.5 from Theorem 1.2 seems rather elementary in algebraic geometry, for the reader’s convenience we provide a short proof in §D of the appendix that does not require any background in the subject.
Remark (Absence of radical expression in general).
For $p,q\in\mathbb{Q}$, the minimal polynomial of $r_{p,q}(\mathbb{Z}_{3})$ may be written explicitly. Indeed, in view of Theorem 1.2 or the equivalent equation-system (4.3) concerning directly $r_{p,q}(\mathbb{Z}_{3})$, one can use the standard method of resultant (see the proof of Corollary 1.5 in §D of the appendix for a quick explanation) to derive an integer-coefficient polynomial (and then an irreducible one) having $r_{p,q}(\mathbb{Z}_{3})$ as a root. For instance, using computer algebra systems (e.g., Mathematica, Matlab), we obtain the minimal polynomial of $r_{3,6}(\mathbb{Z}_{3})$:
@@
1.2. Optimal hypercontractive constants for biased Bernoulli random variables
For the study of hypercontractive phenomena on the discrete probability space $(\Omega,\pi)$, it is convenient to use the notation of hypercontractive random variables introduced by Krakowiak and Szulga (see, e.g., [KS88], [KS91] and [O’D14, Definition 9.13]). For $1<p<q<\infty$ and $0\leq r<1$, a real-valued random variable $X\in L^{q}$ is said to be $(p,q,r)$-hypercontractive if
$\|1+r\rho X\|_{q}\leq\|1+\rho X\|_{p},\qquad\forall\rho\in\mathbb{R}.$
The largest $r$ for which the $(p,q,r)$-hypercontractive inequality holds is called the optimal hypercontractivity constant of the random variable $X$, and is denoted by $\sigma_{p,q}(X)$.
Let $\lambda\in(0,1/2]$. The biased Bernoulli random variable $X_{\lambda}$ is defined by
(1.9) $\mathbb{P}(X_{\lambda}=1-\lambda)=1-\mathbb{P}(X_{\lambda}=-\lambda)=\lambda.$
The optimal hypercontractivity constant $\sigma_{p,q}(X_{\lambda})$ is, in this case, simply denoted by $\sigma_{p,q}(\lambda)$. The $\lambda=1/2$ case is the classical symmetric two-point hypercontractivity inequality, i.e.,
$\sigma_{p,q}(1/2)=r_{p,q}(\mathbb{Z}_{2})=\sqrt{(p-1)/(q-1)}.$
For $\lambda\not=1/2$, the constant $\sigma_{p,q}\left(\lambda\right)$ is called the asymmetric (two-point) hypercontractive constant.
Determining the asymmetric hypercontractive constant $\sigma_{p,q}(\lambda)$ is a highly nontrivial task, even in specific cases. Talagrand [Tal94] established the first nontrivial estimate. Oleszkiewicz [Ole03] derived the first exact formulas when $p=2$ (or $q=2$). Wolff not only extended the Oleszkiewicz formula to general discrete spaces but also obtained analogous results, in a slightly less precise form, for all $2\leq p<q$ and all $1<p<q\leq 2$. Specifically, Wolff [Wol07, Theorems 2.1 and 2.2] showed that $\sigma_{p,q}(\lambda)$ is equivalent (up to a universal numerical constant) to an explicit constant $\widetilde{\sigma}_{p,q}(\lambda)$. He [Wol07, Theorem 3.1] also proved that the hypercontractivity constant of any finite discrete measure is bounded by $\sigma_{p,q}(\lambda)$, where $\lambda$ denotes the mass of the measure’s smallest atom. In particular, by adapting Wolff’s argument, one may show that for all $1<p<q<\infty$,
$r_{p,q}(\mathbb{Z}_{3})=\sigma_{p,q}(1/3).$
To the best of our knowledge, the exact Oleszkiewicz formula for all $1<p<q<\infty$ remains open. We now state our second main result. Using ideas of proving Theorem 1.2, we have
Theorem 1.6.
For $\lambda\in(0,1/2)$ and $1<p<q<\infty$, the hypercontractive constant for the $\lambda$-biased Bernoulli random variable satisfies
$\sigma_{p,q}(\lambda)=\frac{(1-y)\left(x+\frac{\lambda}{1-\lambda}\right)}{(1-x)\left(y+\frac{\lambda}{1-\lambda}\right)},$
with $(x,y)$ is the solution in $(0,1)^{2}$ of the system
(1.10) $\left\{\begin{aligned} &\frac{\big(\lambda+(1-\lambda)x^{p}\big)^{\frac{1}{p}}}{\lambda+(1-\lambda)x}=\frac{\big(\lambda+(1-\lambda)y^{q}\big)^{\frac{1}{q}}}{\lambda+(1-\lambda)y},\\ &\frac{(1-x)(1-x^{p-1})}{\lambda+(1-\lambda)x^{p}}=\frac{(1-y)(1-y^{q-1})}{\lambda+(1-\lambda)y^{q}}.\end{aligned}\right$.
In the particular case $q=p^{*}$, one has
(1.11) $\sigma_{p,p^{*}}(\lambda)=\frac{\sinh\left(\frac{1}{p^{*}}\log\frac{1-\lambda}{\lambda}\right)}{\sinh\left(\frac{1}{p}\log\frac{1-\lambda}{\lambda}\right)}.$